A small source of sound vibrating at frequency 500 Hz is rotated in a circle of radius 100/π cm at a constant angular speed of 5.0 revolutions per second. A listener situates himself in the plane of the circle. Find the minimum and the maximum frequency of the sound observed. Speed of sound in air = 382 m s-1.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. Here given f base s = 16 × 10^3 Hz Apparent frequency f’ = 20 × 10^3 Hz (greater than that value) Let the velocity of the observer = v base o Given v base s = 0 So 20 * 10^3 = (330 + V base o/330 + 0) */ 16 * 10^3 ⇒ (330 + V base o) = 20 * 330/16 ⇒ V base o = 20 * 330 – 16 * 330/4 = 330/4 m/s = 297 km/h b) This speed is not practically attainable ordinary cars.